OpenStudy (anonymous):

Use the limit definition of derivative to find: 3x^2-2x Show me steps if you can. I just need an answer to confirm that my answer was right.

OpenStudy (lgbasallote):

remind me what the definition was...is it \[\huge \lim_{h \rightarrow 0} \frac{f(x+h) - f(x)}h\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[f(x+h)=3(x+h)^2-2(x+h)=3(x^2+2xh+h^2)-2x-2h\] \[=3x^2+6xh+3h^2-2x-2h\] is a start

OpenStudy (anonymous):

nah, I actually done all the steps...just need someone to confirm my answer for me.

OpenStudy (anonymous):

from this you subtract \(x^2-2x\) so every term without an \(h\) in it is gone then factor out the \(h\) and cancel, or else just divide each term by \(h\) since there is an \(h\) in each one

OpenStudy (anonymous):

answer is \(6x-2\) next week you will do this in your head in half a second at most

OpenStudy (anonymous):

oh, thanks...I was right.

OpenStudy (anonymous):

yw